OpenStudy (scienceteen8):

PLEASE HELP I would give a medal and will fan Of the students in Rosalie’s dance class, 29 out of 59 have participated in dance class before. About how many of her students have participated in dance class before? A.0 B.1/2 C.1 D.2

OpenStudy (anonymous):

Since it said "ABOUT how many", it's looking for an approximation. Out of all of the options, what seems closest to \[\large \sf \frac{29}{59}\] ?

OpenStudy (scienceteen8):

ok

OpenStudy (anonymous):

Well, which one?

OpenStudy (scienceteen8):

D

OpenStudy (anonymous):

\[\large \sf \frac{29}{59}\] is closest to 2? Are you sure?

OpenStudy (scienceteen8):

no im not sure lol

OpenStudy (anonymous):

Here, I'll give you a little hint. \[\large \sf \frac{29+1}{59+1}\]

OpenStudy (scienceteen8):

1 or 1/2?

OpenStudy (scienceteen8):

i think 1/2

OpenStudy (anonymous):

Yes, it would be 1/2. When the top number is about half of the bottom number, it will be closest to 1/2. If the top number is close to the bottom number, it would be closer to 1.

OpenStudy (scienceteen8):

ok thnx

OpenStudy (anonymous):

np